# Extending class group action attacks via sesquilinear pairings

### v1

@@ Abstract
We introduce a new tool for the study of isogeny-based cryptography, namely pairings which are sesquilinear (conjugate linear) with respect to the $\mathcal{O}$-module structure of an elliptic curve with CM by an imaginary quadratic order $\mathcal{O}$. We use these pairings to study the security of problems based on the class group action on collections of oriented ordinary or supersingular elliptic curves. This extends work of [4] and [13].
@@
Isogeny-based cryptography Pairings Elliptic Curves.
@@ 1 Introduction
The use of isogeny graphs in cryptography dates to [8, 10, 24]. The latter proposals were for public-key cryptography based on an ordinary isogeny graph. In particular, the class group $\operatorname{Cl}(\mathcal{O})$ of an order $\mathcal{O}$ in an imaginary quadratic field $K$ acts on the set of ordinary elliptic curves over $\overline{\mathbb{F}}_{p}$ with CM by $\mathcal{O}$. For efficiency, CSIDH was proposed [6], making use of supersingular curves with an action by the class group of the Frobenius field. More recently, this was generalized to OSDIH [9], making use of other imaginary quadratic fields in the endomorphism algebra. This paper concerns oriented elliptic curves, which refers to attaching the data of an embedding of a particular imaginary quadratic order $\mathcal{O}$ into the endomorphism ring. All these public-key proposals are examples of class group actions on oriented curves.
The security of these schemes relies on variants of the Diffie-Hellman problem for the class group action. The security of these problems has drawn a great deal of interest, and not all instances of the problem have so far proven to be secure. If the class group is even, the decisional Diffie-Hellman problem is broken by the use of genus theory [7, 5]. These papers make use of the Weil and Tate pairings to compute certain associated characters. More recently, [4] makes use of generalizations of Weil and Tate pairings to break certain instances of the class group action problem (i.e., determining which class group element takes one given oriented curve to another) when the discriminant has a large smooth square factor and the degree is known. Pairings have also appeared in the study of oriented elliptic curves in [17], to navigate the isogeny graph.
The attacks in [4] use pairings to reduce a hidden isogeny problem with known degree for the class group action to the SIDH problem recently broken using higher dimensional abelian varieties [2, 20, 23]. In short, if the degree of a secret isogeny $\phi:E\rightarrow E^{\prime}$ is known, and it is known that $\phi P\in\mathbb{Z}P^{\prime}$ for $P\in E$ and $P^{\prime}\in E^{\prime}$, then we can make use of a relationship of the form
@@
by solving a discrete logarithm to obtain the relationship $k^{2}\equiv\deg\phi\pmod{m}$, and thereby solve for $k$. With this, we (essentially) obtain the image $\phi P$ of $P$, which is the type of information provided in the SIDH problem. The classical SIDH problem (for which we now have efficient methods) requires the image of two basis points, and this provides only one. To close the gap, [4] uses results of [13] which reduce SIDH1, in which only one torsion point is provided, to classical SIDH, provided the order of the point is square. More recent work presented but not yet available [3] uses pairings to generalize the SIDH attacks so that torsion images of any sufficiently large subgroup suffice.
These attacks require that the degree of the secret isogeny is known, which is the case in some implementations; see [4] for some details. Furthermore, in [13, Lemma 14], the authors give a heuristic reduction from the group action problem to the same problem with known degree. In this paper we will assume throughout that the degree of the secret isogeny is known.
In this paper we introduce a new tool for understanding these results and pushing such attacks further. In [26], certain new generalized pairings $\widehat{W}$ and $\widehat{T}$ (generalizing the usual Weil and Tate pairings) are defined, which are $\mathcal{O}$-sesquilinear, meaning that
@@
In particular, we can now assume only that $\phi P\in\mathcal{O}P^{\prime}$ and obtain a relationship
@@
The paper [4] provides a taxonomy of known generalized pairings, but all of these are only $\mathbb{Z}$-bilinear with image in $\mathbb{F}_{q}^{*}$.
One important difference of these sesquilinear pairings from the generalized pairings previously considered is their non-degeneracy. In [4], there is a classification theorem for cyclic self-pairings, namely homogeneous degree-$2$ functions $f_{m}:C\rightarrow\mu_{m}$ where $C$ is a cyclic subgroup of $E[m]$ whose image under $f_{m}$ spans $\mu_{m}$. As shown in [4], these can often be formed from generalizations of the classical reduced $m$-Tate pairing; for a simple such construction, see [4, Example 4.4]. In particular, such functions are helpful when they are compatible with oriented isogenies $\phi$, meaning $f_{m}\circ\phi=f_{m}^{\deg\phi}$. They essentially show that compatible cyclic self-pairings can only be non-trivial for $m$ dividing the discriminant $\Delta_{\mathcal{O}}$ of $\mathcal{O}$. This limits the applicability of their attacks on the class group action to situations where the discriminant has a good factorization. We demonstrate that by extending to $\mathcal{O}$-sesquilinear pairings, whose domain is not $\mathbb{Z}$-cyclic but instead $\mathcal{O}$-cyclic, we obtain many more non-trivial self-pairings to work with.
The use of these new $\mathcal{O}$-sesquilinear pairings offers several clarifying conceptual advantages, and partially answers several of the open problems posed in [4]. However, they are not a magic bullet: we show (Theorem 4.4) that the computation of these pairings is essentially equivalent to the computation of the $\mathcal{O}$-orientation, provided discrete logarithms are efficient in $\mu_{m}$ (for example, if $m$ is smooth).
@@
    We introduce the new $\mathcal{O}$-sesquilinear pairings $\widehat{W}$ and $\widehat{T}$ in the cryptographic context.
@@
    We show an equivalence between computation of an $\mathcal{O}$-orientation and the computation of $\mathcal{O}$-sesquilinear pairings for $m$ smooth (Theorem 4.4).
@@
    Corollary 1 and Theorem 6.1 (described in more detail below) provide evidence for a trade-off between the amount of known level structure of a secret isogeny $\phi:E\to E^{\prime}$ of degree $d$ and how much of the endomorphism rings of $E$ and $E^{\prime}$ we need to represent to find $\phi$. As shown in [27], the fixed-degree isogeny problem with full level structure is equivalent to finding a representation of the full endomorphism ring of $E$ and $E^{\prime}$, while [2, 20, 23] show that the fixed-degree isogeny problem with minimal level structure requires no knowledge of even a partial representation of the endomorphism rings of $E$ and $E^{\prime}$. As described in Cryptographic contributions items 2 and 3 below, knowledge of an intermediate level structure can be combined with a representation of only “half” of the endomorphism rings of $E$ and $E^{\prime}$, to provide attacks on hidden isogenies of known degree. See also the work in [13], which explores varying amounts of level structure.
@@
    We extend the applicability of the (sometimes polynomial) attacks from [4] on the class group action problem (Section 8). These attacks run for smooth $m$ dividing the discriminant. We recover these attacks using the new pairings in a slightly different way, with the advantage that our pairing computations do not require going to a large field extension. This partially addresses one of the open questions of [4, Section 7]. Example 3 gives an explicit situation in which the reach of attacks is extended.
@@
    We demonstrate a pairing-based reduction from SIDH1 to SIDH in the oriented situation for $E[m]$, where $m$ is smooth and coprime to the discriminant (Theorem 6.1), resulting in an attack when $m^{2}>\deg\phi$. This partially addresses the first and second open problems in [4, Section 7]. Existing attacks on SIDH1 (which apply without orientation information) require $m>\deg\phi$.
@@
    We reduce the hard problem underlying FESTA [1] to finding an orientation of the secret isogeny $\phi:E\rightarrow E^{\prime}$ (i.e. an orientation of both curves and the isogeny between them) (Corollary 1). This follows from an attack on the Diagonal SIDH Problem (Theorem 7.1).
@@
    We show how these pairings, using orientation information, easily reveal partial information on the image of a torsion point $P$ of order $m$ for $m$ smooth (Theorem 5.1). This results in an algorithm to break CSIDH or other class-group-based schemes in square root of exponential time by running the SIDH attack on $\sqrt{m}$ candidate $m$-torsion points as images under $\phi$ (Remark 4).
@@
    Our results should be considered a cautionary tale for the design of decisional problems based on torsion point images, such as in [21], since the possible images of torsion points is restricted. We discuss this in Remark 5.
@@
    In the supersingular case, we demonstrate a method of finding the secret isogeny in the presence of two independent known orientations (which amounts to an explicit subring of the endomorphism ring of rank $4$), provided the secret isogeny is oriented for both orientations. This is not a surprise, as this problem could be solved by the KLPT algorithm if the endomorphisms are obtained by walking the graph (see [12], and also [18, 27]), but it provides a new method via a simple reduction to the SIDH problem. (Section 9.)
@@ 2 Background
We study elliptic curves, typically denoted $E$, $E^{\prime}$ etc., defined over finite fields, denoted by $\mathbb{F}$ in general. Denote an algebraic closure of $\mathbb{F}$ by $\bar{\mathbb{F}}$. The identity on $E$ is denoted $\mathcal{O}_{E}$, and $\operatorname{End}(E)$ is the endomorphism ring over $\overline{\mathbb{F}}$. We study imaginary quadratic fields, denoted $K$ in general, and orders in such fields, denoted by $\mathcal{O}$, $\mathcal{O}^{\prime}$ etc. Greek letters typically denote elements of the orders. When considering the action of an element $\alpha\in\mathcal{O}$ on a point $P$, we write $[\alpha]P$. The Greek letter $\phi$ always refers to an isogeny. For ease of notation, we write $\phi P$ instead of $\phi(P)$. Throughout the paper, we write $\mu_{m}$ for the copy of $\mu_{m}$ in a finite field.
@@
We study $\mathcal{O}$-oriented elliptic curves over finite fields, which are curves together with the information of an embedding $\iota:\mathcal{O}\rightarrow\operatorname{End}(E)$. This extends to an embedding of the same name, $\iota:K\rightarrow\mathbb{Q}\otimes_{\mathbb{Z}}\operatorname{End}(E)$, and the $\mathcal{O}$-orientation is called primitive if $\iota(K)\cap\operatorname{End}(E)=\mathcal{O}$. If the index $[\iota(K)\cap\operatorname{End}(E):\iota(\mathcal{O})]$ is coprime to $n$, we say the orientation is $n$-primitive. Given an $\mathcal{O}$-orientation, there is a unique $\mathcal{O}^{\prime}\supseteq\mathcal{O}$ for which $\iota$ becomes a $\mathcal{O}^{\prime}$-primitive orientation, namely $\mathcal{O}^{\prime}=\iota(K)\cap\operatorname{End}(E)$. Given an elliptic curve $E$ with an $\mathcal{O}$-orientation, we define the relative conductor of $\mathcal{O}$ to be the index $[\mathcal{O}^{\prime}:\mathcal{O}]$, for which the orientation is $\mathcal{O}^{\prime}$-primitive.

### v2

@@ Abstract
We introduce a new tool for the study of isogeny-based cryptography, namely pairings which are sesquilinear (conjugate linear) with respect to the $\mathcal{O}$-module structure of an elliptic curve with CM by an imaginary quadratic field $\mathcal{O}$. We use these pairings to study the security of problems based on the class group action on collections of oriented ordinary or supersingular elliptic curves. This extends work of [CHM+23] and [FFP24].
@@
Isogeny-bas/Users/josephmacula/Desktop/untitled foldered cryptography Pairings Elliptic Curves.
@@ 1 Introduction
The use of isogeny graphs in cryptography dates to [CLG09, Cou06, RS06]. The latter proposals were for public-key cryptography based on an ordinary isogeny graph. In particular, the class group $\operatorname{Cl}(\mathcal{O})$ of an order $\mathcal{O}$ in an imaginary quadratic field $K$ acts on the set of ordinary elliptic curves over $\overline{\mathbb{F}}_{p}$ with CM by $\mathcal{O}$. For efficiency, CSIDH was proposed [CLM+18], making use of supersingular curves with an action by the class group of the Frobenius field. More recently, this was generalized to OSDIH [CK20], making use of other imaginary quadratic fields in the endomorphism algebra. Recently, SIDH adaptations based on related ideas have been proposed [BF23]. Our paper concerns oriented elliptic curves, which refers to attaching the data of an embedding of a particular imaginary quadratic order $\mathcal{O}$ into the endomorphism ring. All these public-key proposals are examples of class group actions on oriented curves.
The security of these schemes relies on variants of the Diffie-Hellman problem for the class group action. The security of these problems has drawn a great deal of interest, and not all instances of the problem have so far proven to be secure. If the class group is even, the decisional Diffie-Hellman problem is broken by the use of genus theory [CSV22, CHVW22]. These papers make use of the Weil and Tate pairings to compute certain associated characters. More recently, [CHM+23] makes use of generalizations of Weil and Tate pairings to break certain instances of the class group action problem (i.e., determining which class group element takes one given oriented curve to another) when the discriminant has a large smooth square factor and the degree is known. Pairings have also appeared in the study of oriented elliptic curves in [IJ13], to navigate the isogeny graph. For other interactions between pairings and isogeny-based cryptography, see [KT19, Rei23].
The attacks in [CHM+23] use pairings to reduce a hidden isogeny problem with known degree for the class group action to the SIDH problem recently broken using higher dimensional abelian varieties [CD23, MMP+23, Rob23b]. In short, if the degree of a secret isogeny $\phi:E\rightarrow E^{\prime}$ is known, and it is known that $\phi P\in\mathbb{Z}P^{\prime}$ for $P\in E$ and $P^{\prime}\in E^{\prime}$, then we can make use of a relationship of the form
@@
by solving a discrete logarithm to obtain the relationship $k^{2}\equiv\deg\phi\pmod{m}$, and thereby solve for $k$. With this, we (essentially) obtain the image $\phi P$ of $P$, which is the type of information provided in the SIDH problem. The classical SIDH problem (for which we now have efficient methods) requires the image of two basis points, and this provides only one. To close the gap, [CHM+23] uses results of [FFP24] which reduce SIDH1, in which the image of only one torsion point is provided, to classical SIDH, provided the order of the point is square. More recent work presented but not yet available [CDM+24] uses pairings to generalize the SIDH attacks so that torsion images of any sufficiently large subgroup suffice.
These attacks require that the degree of the secret isogeny is known. This is the case in constant-time implementations aimed at preventing side-channel attacks such as those in [CVCCD+19]; see [CHM+23] for more details. Furthermore, in [FFP24, Lemma 14], the authors give a heuristic reduction from the group action problem to the same problem with known degree. In this paper we will assume throughout that the degree of the secret isogeny is known.
In this paper we introduce a new tool for understanding these results and pushing such attacks further. In [Sta24], certain new generalized pairings $\widehat{W}$ and $\widehat{T}$ (generalizing the usual Weil and Tate pairings) are defined, which are $\mathcal{O}$-sesquilinear, meaning that
@@
In particular, we need now assume only that $\phi P\in\mathcal{O}P^{\prime}$ and obtain a relationship
@@
The paper [CHM+23] provides a taxonomy of known generalized pairings, but all of these are only $\mathbb{Z}$-bilinear with image in $\mathbb{F}_{q}^{*}$.
One important difference of these sesquilinear pairings from the generalized pairings previously considered is their non-degeneracy. In [CHM+23], there is a classification theorem for cyclic self-pairings compatible with oriented endomorphisms. These are functions $f_{m}:C\rightarrow\mu_{m}$ where $C$ is a cyclic subgroup of $E[m]$ whose image under $f_{m}$ spans $\mu_{m}$, with the following properties: $f(\lambda P)=f(P)^{\lambda^{2}}$, $\iota(\sigma)(P)\in C$, and $f(\iota(\sigma)P)=f(P)^{N(\iota(\sigma))}$ for $\iota$ an orientation of a given imaginary quadratic order $\mathcal{O}$, $\sigma\in\mathcal{O}$, and $P\in C$. They essentially show that such pairings can only be non-trivial for $m$ dividing the discriminant $\Delta_{\mathcal{O}}$ of $\mathcal{O}$.
The requirement that $m$ divide $\Delta_{\mathcal{O}}$ limits the applicability of their attacks on the class group action to situations where the discriminant has a good factorization. We demonstrate that by extending to $\mathcal{O}$-sesquilinear pairings, whose domain is not $\mathbb{Z}$-cyclic but instead $\mathcal{O}$-cyclic, we obtain many more non-trivial self-pairings to work with.
The use of these new $\mathcal{O}$-sesquilinear pairings offers several clarifying conceptual advantages, and partially answers several of the open problems posed in [CHM+23]. However, they are not a magic bullet: we show (Theorem 4.4) that the computation of these pairings is essentially equivalent to the computation of the $\mathcal{O}$-orientation, provided discrete logarithms are efficient in $\mu_{m}$ (for example, if $m$ is smooth).
@@
    We introduce the new $\mathcal{O}$-sesquilinear pairing $\widehat{T}$ in the cryptographic context.
@@
    We show an equivalence between computation of an $\mathcal{O}$-orientation and the computation of $\mathcal{O}$-sesquilinear pairings for nice $m$ (Theorem 4.4).
@@
    Corollary 1 and Theorem 6.1 (described in more detail below) provide evidence for a trade-off between the amount of known level structure of a secret isogeny $\phi:E\to E^{\prime}$ of degree $d$ and how much of the endomorphism rings of $E$ and $E^{\prime}$ we need to represent to find $\phi$. As shown in [Wes22], the fixed-degree isogeny problem with full level structure is equivalent to finding a representation of the full endomorphism ring of $E$ and $E^{\prime}$, while [CD23, MMP+23, Rob23b] show that the fixed-degree isogeny problem with minimal level structure requires no knowledge of even a partial representation of the endomorphism rings of $E$ and $E^{\prime}$. As described in Cryptographic contributions items 2 and 3 below, knowledge of an intermediate level structure can be combined with a representation of only “half” of the endomorphism rings of $E$ and $E^{\prime}$, to provide attacks on hidden isogenies of known degree. See also the work in [FFP24], which explores varying amounts of level structure.
@@
    We extend the applicability of the (sometimes polynomial) attacks from [CHM+23] on the class group action problem (Section 8). These attacks run for smooth $m$ dividing the discriminant. We recover these attacks using the new pairings in a slightly different way, with the advantage that our pairing computations do not require going to a large field extension. This partially addresses one of the open questions of [CHM+23, Section 7]. Example 3 gives an explicit situation in which the reach of polynomial attacks is strictly extended.
@@
    We demonstrate a pairing-based reduction from SIDH1 to SIDH in the oriented situation for $E[m]$, where $m$ is smooth and coprime to the discriminant (Theorem 6.1), resulting in an attack when $m^{2}>\deg\phi$. This partially addresses the first and second open problems in [CHM+23, Section 7]. Existing attacks on SIDH1 (which apply without orientation information) require $m>\deg\phi$.
@@
    We reduce the hard problem underlying FESTA [BMP23] to finding an orientation of the secret isogeny $\phi:E\rightarrow E^{\prime}$ (i.e. an orientation of both curves and the isogeny between them) (Corollary 1). This follows from an attack on the Diagonal SIDH Problem (Theorem 7.1).
@@
    We show how these pairings, using orientation information, easily reveal partial information on the image of a torsion point $P$ of order $m$ for $m$ smooth (Theorem 5.1). This results in an algorithm to break class-group-based schemes by running the SIDH attack on $\sqrt{\deg(\phi)}$ candidate torsion points as images under $\phi$ (Remark 5).
@@
    Our results should be considered a cautionary tale for the design of decisional problems based on torsion point images, such as in [MOT20], since the possible images of torsion points is restricted. We discuss this in Remark 6.
@@
    In the supersingular case, we demonstrate a method of finding the secret isogeny in the presence of two independent known orientations (which amounts to an explicit subring of the endomorphism ring of rank $4$), provided the secret isogeny is oriented for both orientations. This is not a surprise, as this problem could be solved by the KLPT algorithm if the endomorphisms are obtained by walking the graph (see [EHL+20], and also [KLPT14, Wes22]), but it provides a new method via a simple reduction to the SIDH problem. (Section 9.)
@@ 2 Background
We study elliptic curves, typically denoted $E$, $E^{\prime}$ etc., defined over finite fields, denoted by $\mathbb{F}$ in general. Denote an algebraic closure of $\mathbb{F}$ by $\bar{\mathbb{F}}$. The identity on $E$ is denoted $\infty$, and $\operatorname{End}(E)$ is the endomorphism ring over $\overline{\mathbb{F}}$. We study imaginary quadratic fields, denoted $K$ in general, and orders in such fields, denoted by $\mathcal{O}$, $\mathcal{O}^{\prime}$ etc. Greek letters typically denote elements of the orders. We denote the norm of an element $\lambda$ of a given order by $N(\lambda)$. When considering the action of an element $\alpha\in\mathcal{O}$ on a point $P$, we write $[\alpha]P$. The Greek letter $\phi$ always refers to an isogeny and $\widehat{\phi}$ always denotes the dual isogeny of $\phi$. For ease of notation, we write $\phi P$ instead of $\phi(P)$. Throughout the paper, we write $\mu_{m}$ for the copy of $\mu_{m}$ in a finite field.
@@
We study $\mathcal{O}$-oriented elliptic curves over finite fields, which are curves together with the information of an embedding $\iota:\mathcal{O}\rightarrow\operatorname{End}(E)$. This extends to an embedding of the same name, $\iota:K\rightarrow\mathbb{Q}\otimes_{\mathbb{Z}}\operatorname{End}(E)$, and the $\mathcal{O}$-orientation is called primitive if $\iota(K)\cap\operatorname{End}(E)=\iota(\mathcal{O})$. If the index $[\iota(K)\cap\operatorname{End}(E):\iota(\mathcal{O})]$ is coprime to $n$, we say the orientation is $n$-primitive. Given an $\mathcal{O}$-orientation, there is a unique $\mathcal{O}^{\prime}\supseteq\mathcal{O}$ for which $\iota$ becomes a $\mathcal{O}^{\prime}$-primitive orientation, namely $\iota(\mathcal{O}^{\prime})=\iota(K)\cap\operatorname{End}(E)$. Given an elliptic curve $E$ with an $\mathcal{O}$-orientation, we define the relative conductor of $\mathcal{O}$ to be the index $[\mathcal{O}^{\prime}:\mathcal{O}]$, for which the orientation is $\mathcal{O}^{\prime}$-primitive.
@@
2.3 Cyclic self-pairings
CSIDH, introduced in [CLM+18], relies for its security on the presumed hardness of the following instance of the vectorization problem: given a (large) prime $p\equiv 3\pmod{4}$ and two supersingular curves $E$ and $E^{\prime}$ over $\mathbb{F}_{p}$, find the ideal class $[\mathfrak{a}]$ of $\mathcal{O}=\mathbb{Z}[\sqrt{-p}]$ such that $E^{\prime}=[\mathfrak{a}]E$. More generally, the vectorization problem can be phrased as follows: given two supersingular elliptic curves $E,E^{\prime}$ primitively oriented by an imaginary quadratic order $\mathcal{O}$ and known to be connected by the action of the ideal class group cl$(\mathcal{O})$ of $\mathcal{O}$, find $[\mathfrak{a}]\in\text{cl}(\mathcal{O})$ such that $E^{\prime}=[\mathfrak{a}]E$. This is also known as a class-group-action problem.
This paper builds on the previous work of [CHM+23]. There, the authors ask whether the attack on SIDH [CD23, MMP+23, Rob23b] that renders the protocol insecure can be applied to solving the vectorization problem. In brief, they note that one can treat the SIDH attack as an oracle, which when given the degree $d$ of a secret $\mathbb{F}_{q}$-rational isogeny $\phi$ between curves $E$ and $E^{\prime}$ defined over $\mathbb{F}_{q}$ and knowledge of its action on a basis of $E[m]$ for $m$ coprime to $d$ and $m^{2}>4d$, returns $\phi$. Assuming the degree $d$ of $\phi$ is known, the question of whether this oracle can answer the vectorization problem therefore reduces to the question of whether one can determine the action of $\phi$ on a basis of $E[m]$ for suitable $m$.
The following example, reproduced directly from [CHM+23], is instructive. Assume the context of CSIDH, i.e., that the relevant order is $\mathbb{Z}[\sqrt{-p}]$. Choosing $m$ to be the power of a small prime $l$ coprime to $d$ that splits in $\mathbb{Q}(\sqrt{-p})$, $E[m]$ has a basis $\{P,Q\}$ consisting of eigenvectors of the Frobenius endomorphism $\pi_{p}$. Since by assumption the curves $E,E^{\prime}$ and the isogeny $\phi$ are all defined over $\mathbb{F}_{p}$, $E^{\prime}[m]$ has a basis $\{P^{\prime},Q^{\prime}\}$ consisting of eigenvectors of $\pi_{p}$ and $\phi P=rP^{\prime}$, $\phi Q=sQ^{\prime}$ for $r,s\in\left(\mathbb{Z}/m\mathbb{Z}\right)^{\times}$. The bilinearity and compatiblity with isogenies of the $m$-Weil pairing imply that
$e_{m}(P^{\prime},Q^{\prime})=e_{m}(P,Q)^{rsd}$ (1)
With Miller’s algorithm [Mil04], computation of the $m$-Weil pairing is efficient. Since $m$ is a power of a small prime, also efficient is computation of discrete logarithms. Thus, given knowledge of $d$ it remains to determine one of $r$ or $s$. Yet properties of the $m$-Weil pairing imply that $e_{m}(P^{\prime},P^{\prime})=1$; thus, one cannot compute $e_{m}(P^{\prime},P^{\prime})=e_{m}(P,P)^{r^{2}d}$ to find $r$.
In [CHM+23], this obstacle is surmounted via the construction of cyclic self-pairings that are compatible with $\mathcal{O}$-oriented isogenies. A cyclic self-pairing is a function $f$ defined on a finite cyclic subgroup $C$ of an elliptic curve $E/\mathbb{F}$ with the property that
$f(rP)=f(P)^{r^{2}}\hskip 7.11317pt\text{for all }P\in C\text{ and }r\in\mathbb{Z}.$
When $E$ and $E^{\prime}$ are two curves over $\mathbb{F}$ with orientations of an imaginary quadratic order $\mathcal{O}$ by $\iota,\iota^{\prime}$, respectively, two self-pairings $f$ and $f^{\prime}$ defined on finite subgroups $C$ of $E$ and $C^{\prime}$ of $E^{\prime}$ are compatible with $\mathcal{O}$-oriented isogenies $\phi:E\to E^{\prime}$ when $\phi(C)\subset C^{\prime}$ and $f^{\prime}(\phi P)=f(P)^{\deg\phi}$. If $\phi$ is an $\mathcal{O}$-oriented isogeny from $E$ to $E^{\prime}$ of degree $d$ coprime to an integer $m$ such that $E$ and $E^{\prime}$ have non-trivial cyclic self-pairings $f$ and $f^{\prime}$ compatible with $\mathcal{O}$-oriented isogenies on cyclic subgroups $\langle P\rangle$, $\langle P^{\prime}\rangle$, then
$f^{\prime}(P^{\prime})=f(P)^{dr^{2}}$
for some $r\in\mathbb{Z}/m\mathbb{Z}$. If furthermore discrete logarithms are efficiently computable modulo $m$, then the non-triviality of $f$ and $f^{\prime}$ implies that one can efficiently determine $r^{2}$ modulo $m$. Assuming $m$ has a nice factorization, this leaves only a few possibilities for $r$, and one simply guesses by direct computation which one is correct.
The non-trivial self-pairings constructed in [CHM+23] are generalizations on the classical reduced $m$-Tate pairing. We refer the reader to Section 5 of [CHM+23] for further details. Crucially, the order $m$ of non-trivial cyclic self-pairings compatible with $\mathcal{O}$-oriented isogenies must divide $\Delta_{\mathcal{O}}$ ([CHM+23], Proposition 4.8). Furthermore, the existence of such a self-pairing only yields knowledge of the image of a single torsion point $P$ under $\phi$. In [CHM+23], this latter issue is addressed by assuming $m^{2}\mid\Delta_{\mathcal{O}}$. Then one obtains the image of an order $m^{2}$ point $P$ under $\phi$. The authors briefly describe how one can then obtain from this data knowledge of the action of $\phi$ on a basis for $E[m]$. A more systematic description of this reduction in the language of level structure is in Section 5 of [FFP24] (in particular, see corollary 12).
Thus, there are two significant limitations to the scope of this attack. First, $\Delta_{\mathcal{O}}$ must contain a large smooth square factor. Second, in general one must work over a field where the $m^{2}$-torsion is fully rational. In the worst case, this requires a base change to an extension of potentially large degree.
The first of these limitations is addressed in work in preparation by Castryck et. al. [CDM+24], the authors show that with knowledge of the image of $\phi:E\to E^{\prime}$ on a generating set $S$ for a subgroup $G$ with #$\#G>4d$, there is an algorithm to determine $\phi$ (in the sense of computing arbitrary images) that is polynomial in
  1. (1)
    the size of the $S$ and $\log q$, where $q$ is the size of the field over which $E$ and $E^{\prime}$ are defined;
  2. (2)
    the size of the largest prime factor of #$\#G$;
  3. (3)
    the largest degree of the fields of definition of $E[\ell^{\lfloor e/2\rfloor}]$, taken over all prime powers $l^{e}$ dividing #$\#G$.
In particular, this result obviates the need for $\Delta_{\mathcal{O}}$ to contain a smooth square factor; instead, one only needs a smooth factor of size greater than $4d$.
2.4 Level Structure
Many isogeny-based protocols require that some torsion-point information be made publicly known. For example, in SIDH, the image under the secret isogeny $\phi$ of a specified basis $\{P,Q\}$ for $E[m]$ (where $m$ is a power of a small prime coprime to the characteristic of the field $\mathbb{F}$ over which $E$ is defined) is known. As discussed in the last section, in CSIDH the image of a basis $\{P,Q\}$ for $E[m]$ (with $m$ again a power of a small prime coprime to the field characteristic, but also coprime to the degree $d$ of the secret isogeny $\phi$) is known, up to multiplication by an element of $\left(\mathbb{Z}/m\mathbb{Z}\right)^{\times}$. Equivalently, the image under $\phi$ of two order $m$ subgroups of $E$ is known. Both types of torsion-point information are examples of level structure that $\phi$ respects.
Definition 1 ([FFP24])
Let $E$ be an elliptic curve over a finite field $\mathbb{F}$ of characteristic $p$ and $m$ be a positive integer coprime to $p$. Let $\Gamma$ be a subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/m\mathbb{Z})$. A $\Gamma$-level structure of level $m$ on $E$ is a $\Gamma$-orbit of a basis of $E[m]$.
Typically in the context of isogeny problems, one is not interested in level structure per se, but in level structure that a given isogeny $\phi$ respects. That is, given curves $E$ and $E^{\prime}$ both with $\Gamma$-level $m$ structures for a fixed $\Gamma$, $\phi$ maps the specified $\Gamma$-orbit for $E[m]$ to the specified $\Gamma$-orbit for $E^{\prime}[m]$.
There has been much attention paid recently to elliptic curves equipped with a particular level structure. Arpin [Arp24] studies the correspondence of Eichler orders in the quaternion algebra $B_{p,\infty}$ with supersingular elliptic curves over $\overline{\mathbb{F}}_{p}$ equipped with Borel level structure—i.e., where $\Gamma=\left(\{\begin{smallmatrix}*&0\\ *&*\end{smallmatrix}\}\right)$—for $m$ squarefree and coprime to $p$. In [CL24], the authors consider the structure of the supersingular isogeny graph with varying level structures and show that many of these graphs remain Ramanujan. Others investigate the actions of generalized ideal class groups on elliptic curves over finite fields with level structure [GPV23, ACKE+24]. In this paper, we are primarily interested in level structure as a unifying framework for understanding the security of various proposals in isogeny-based cryptography. This framework is described in detail in [FFP24]. In particular, those authors make explicit the implicit level structures in several schemes including SIDH, M-SIDH, CSIDH, and FESTA, and prove several security reductions between various level structures.